Julian needs to buy Advil to feel better. He has $12 to spend on medication. Supposes that he can buy Advil in two dosages: 200mg and 400mg. He only cares about getting better and not about the amount of pills he needs to swallow. A 200mg pill costs $0.30 and a 400mg pill costs $1. What will be in Julian's optimal consumption bundle?

Does the answer change if a 200mg pill costs $0.60? If yes, how?
Answer #1

Total Money to be spent on medication = $12

a) 200mg pill costs $0.3 and 400mg pill costs $1

If he buys 2 pills of 200mg it is equal to one 400mg pill

But, Price of 2 pills of 200mg = 2 * $0.3 = $0.6

and Price of one 400mg pill = $1

So, the effective price of 200mg pill is cheaper than the 400mg pill with both being substitute of one another.

Therefore Julian will consume all of 200mg pill with his income and none of 400mg pill

Optimal Bundle :-

200mg pill Demand = 12/0.3 = 40

400mg pill demand = 0

So, Bundle is (40,0)

b) Yes the answer changes

If Now 200mg pill costs $0.6

So, 2 pills of 200mg costs = 2 * $0.6 = $1.2

But One pill of 400mg still costs = $1

Now, The 400mg pill becomes cheaper so he consumes all of it and none of 200mg pill as Julian only cares about being better and not about the number of pills he needs to swallow.

Optimal Demand :-

200mg pill demand = 0

400mg pill demand = 12/1 = 12

So, Optimal Bundle is (0,12)

Hence, YES the answer changes as optimal bundle in first case was (40,0) and now it's (0,12)